OpenStudy (anonymous):

How much time is required for a bicycle to travel a distance of 100 m at an average speed of 2 m/s

OpenStudy (anonymous):

The bicycle will take 1 second to walk 2 meters. So, it will take: x seconds to walk 100. 2x = 100, so we get: x = 50 seconds.

OpenStudy (anonymous):

Or remember that\[Avg speed = \frac{\Delta s}{\Delta t}\]and we want deltaT, so:\[\Delta t = \frac{\Delta S}{Avgspeed} = \frac{100}{2}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem :-)